# Predictive Data Selection: The Data That Predicts Is the Data That Teaches

### v3

@@ 6 Conclusion
In this work, we introduce PreSelect, a novel predictive data selection tailored for language model pretraining that leverages the predictive strength of the data (i.e. how effectively the normalized losses on the data represent the model’s capabilities) to identify high-quality and low-quality data. Unlike traditional heuristic-based filtering or computationally expensive model-based data selection, PreSelect is not only more lightweight, requiring only the fastText scorer at deployment, but also more principled, thereby bypassing strong human heuristics. In addition, by operating at the document level, PreSelect significantly improves granularity and adaptability in data selection. Experimental results across multiple model scales validate the effectiveness of PreSelect, demonstrating substantial improvements over random selection and recent pretraining data selection baselines in both downstream task performance and computational efficiency. Our findings highlight the potential of predictive strength as a reliable indicator for measuring data quality, paving the way for more efficient and scalable pretraining data selection.
@@ A.1 Identified Positive/Negative Data
We then select examples with score = 1 as the positive(good) data and select negative(bad) data gradually from score = 0 as dicussed in § 2.3.
@@ A.2 FastText Training Details
We use the official fastText python library [^3] https://github.com/facebookresearch/fastText for training the classifier, some important hyper-parameters are listed in Table 5 below. Basically we follow the default setting of fastText training, however, since both our positive examples and negative examples are from pretraining data distribution which tend to be diverse, so we increase the epoch number to 5 for better trianing convergence.
@@ A.4 Extend to Pretraining Corpora
We use DataTrove (Penedo et al., 2024c) which is a library for processing data at very large scale. The final filtering process on pretraining corpora does not any GPU resrouces and the filtering process is done on the machine with 4 AMD EPYC 9654 96-Core Processor.
@@ B.2 Pretraining Corpus
Perplexity Correlation 10%(domain-wise) 30B
@@
For RefinedWeb, we use a version [^4] https://data.commoncrawl.org/contrib/datacomp/DCLM-refinedweb/index.html that after heuristic filtering and deduplication following DCLM and randomly sampled a subset of needed number of tokens. We evenly sampled from each global index and local index. For C4, we use the whole dataset, which is about 198B tokens.
@@ C.2 Full Evaluation Results
Task Random Selection Perplexity Filter FineWeb-Edu Perplexity Correlation DCLM PreSelect

### v4

@@ 6 Conclusion
In this work, we introduce PreSelect, a novel predictive data selection tailored for language model pretraining that leverages the predictive strength of the data (i.e. how effectively the normalized losses on the data represent the model’s capabilities) to identify high-quality and low-quality data. By operating at the document level, PreSelect significantly improves granularity and adaptability in data selection. Experimental results across multiple model scales validate the effectiveness of PreSelect, demonstrating substantial improvements over random selection and recent pretraining data selection baselines in both downstream task performance and computational efficiency.
Acknowledgments
This project is partially supported by NSFC Grant 62306177.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Identified Positive/Negative Data
We then select examples with score = 1 as the positive(good) data and select negative(bad) data gradually from score = 0 as dicussed in § 2.3. We have already give discussion about what kind of data are these positive and negative data in § 4.1, here we list more domains (e.g. top-10 domains) these positive/negative documents from in Figure 5 and Figure 6 respectively.
Figure 5: The Top-10 domains where the identified positive documents from, measured by percentage of positive documents inside that domain.
Figure 6: The Top-10 domains where the identified negative documents from, measured by percentage of negative documents inside that domain.
@@ A.2 FastText Training Details
We use the official fastText python libraryhttps://github.com/facebookresearch/fastText for training the classifier, some important hyper-parameters are listed in Table 5 below. Basically we follow the default setting of fastText training, however, since both our positive examples and negative examples are from pretraining data distribution which tend to be diverse, so we increase the epoch number to 5 for better trianing convergence.
@@ A.4 Extend to Pretraining Corpora
We use DataTrove (Penedo et al., 2024c) which is a library for processing data at very large scale. The final filtering process on pretraining corpora does not need any GPU resrouces and the filtering process is done on the machine with 4 AMD EPYC 9654 96-Core Processor.
@@ A.6 Filtering Computation Overhead
A.7 Additional Analysis on Filtering
A.7.1 Comparison with ScalingFilter
ScalingFilter (Li et al., 2024b) uses the perplexity difference between a large model and a small model as a metric to select data, which can be viewed as a method only contrasts losses between two model scales. To show these intermediate level models actually help, we compare with ScalingFilter under our setting, using llama-65B and llama-7B to select data under ScalingFilter while removing the mid-size models to compute the Quality factor. Results of training 1B model on 30B tokens with our data pool (Table 1 setting) are shown in Table 7 below.
Table 7: The averaged performance comparison between ScalingFilter and PreSelect.
Method Average Benchmark Performance
Random 37.2
ScalingFilter 37.6
PreSelect 40.3
From Table 7, we see that after removing the 4 mid-size models, the results drop by 2.7 points on average. Take a further step, we observe that data with a significant perplexity difference is skewed toward shorter texts, with an average character length of 2300, compared to the original corpus distribution’s average of 3500 characters. We further see that the top domains where these data with significant perplexity difference from contains (regator.com, msbusiness.com, www.yourobserver.com, www.qacollections.com, www.stylebistro.com, libn.com, www.businessmanagementdaily.com) where none of typical high-quality domains exist. These evidence suggests that using the perplexity difference between a large model and a small model tends to select easy and short data.
In addition, a large loss difference between the largest model and the smallest model does not necessarily imply a rank preservation among a series of models. On the other hand, a rank matching between losses and downstream performance of a series of models does not necessarily imply a large losses difference between the largest model and the smallest model. For example, we calculate the spearman correlation between the ScalingFilter perplexity ratios and our PreSelect predictive strength, which is 0.0533. And they also have a Pearson correlation of -0.079 which indicate a low correlation between these two metrics (measured based on the actual predictive strength score where we calculate based on the sampled subset), indicating the effectiveness of intermediate level models.
A.7.2 Targeting Specific Downstream Task
Since our setting is pre-training, we didn’t intend to choose a specific task as the target. However, in our initial experiments, we did explore such differences. For example, when choosing HellaSwag to represent intelligence that led to a different model ranking from using the average. This results in a performance improvement on knowledge-intensive tasks, indicated by 5% lower losses on wiki-related raw texts while having much higher (worse) losses on other domains such as 8% on math and 16% on code, at a 400M model and 8B token training scale. Afterwards, we consistently choose the average performance to represent intelligence under pre-training setting, but We think this is evidence that our method actually predicts “downstream abilities” beyond only pre-training scale.
A.7.3 Mitigating Sensitive Contents
In our observation, literature and knowledge related contents are more reflective of downstream tasks (generally have a better rank alignment) which includes essays and some adult contents. We believe filtering adult content is an orthogonal direction that could be done separately, which is not the focus of this work. For example, we think a rule-based pre-processing step to filter adult contents out could greatly mitigate this issue.
@@ B.2 Pretraining Corpus
Perplexity Correlation(DD) 10%(domain-wise) 30B
Perplexity Correlation(DP) 10% 30B
@@
For RefinedWeb, we use a versionhttps://data.commoncrawl.org/contrib/datacomp/DCLM-refinedweb/index.html that after heuristic filtering and deduplication following DCLM and randomly sampled a subset of needed number of tokens. We evenly sampled from each global index and local index. For C4, we use the whole dataset, which is about 198B tokens.
@@ C.2 Full Evaluation Results
Task Random Selection Perplexity Filter FineWeb-Edu Perplexity Correlation(DD) DCLM PreSelect